23. Calculating Costs of Issuing Debt Home Improvement, Inc. needs to raise $2.20 million to finance plant expansion. In discussions with its investment bank, Home Improvement learns that the bankers recommend a debt issue with a gross proceeds of $1,000 per bond and they will charge an underwriter's spread of 8 percent of the gross proceeds. How many bonds will Home Improvement need to sell in order to receive the $2.20 million they need?

  • 2,376

  • 2,392

  • 2,376,000

  • 2,391,304
Answer #1

Solution:

Gross proceeds = $1000 per bond

Underwriters spread =8%

Net proceeds = 1000- (1000*8%) = $920

Funds need to raise = $22,00,000

Number of bonds need to sell = $22,00,000/920= 2391.30 rounded to 2392 bonds

Option B is correct i.e. 2392